OpenStudy (anonymous):

tell wheter function is quadratic explain x -2,-1,0,1,2 y -8,-1,0,1,8

OpenStudy (anonymous):

No, not quadratic. It appears that one way you could produce these y-values is by cubing the x-value. i.e., y=x^3

OpenStudy (anonymous):

y is it not quadratic?

OpenStudy (anonymous):

Quadratics are specific funtions in the the form y=ax^2+bx+c. As you can see, y=dx^3+fx^2+gx+h does not fit this description. Therefore, it is not a quadratic equation. :)